Whats the lenth of a diagonal of a rectangle when the sides are 6ft and 3ft??
Answer by JulietG(1812)

When there is a right angle, a^2 + b^2 = c^2 where a and b are the sides and c is the hypotenuse.
.
In a rectangle, it's two triangles put together.
The equation is:
6^2 + 3^2 = x^2
36 + 9 = x^2
45 = x^2
6.708 = x